OpenStudy (tennistar):

Simplify x squared minus 6 times x plus 9 all over x squared minus 9.

OpenStudy (tennistar):

x minus 3 all over x plus 3. x plus 6 all over x minus 6. x minus 6 all over x minus 3. x plus 3 all over x plus 6.

OpenStudy (anonymous):

\[\frac{x^2 - 6x + 9}{x^2 - 9} = \frac{(x-3)(x-3)}{(x-3)(x+3)} = \frac{x-3}{x+3}\]

OpenStudy (tennistar):

ok, now I get it thanks

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

can u explain some more on how u do that